### v1

# CasTGAN: Cascaded Generative Adversarial Network for Realistic Tabular Data Synthesis

Abstract
Generative adversarial networks (GANs) have drawn considerable attention in recent years for their proven capability in generating synthetic data which can be utilized for multiple purposes. While GANs have demonstrated tremendous successes in producing synthetic data samples that replicate the dynamics of the original datasets, the validity of the synthetic data and the underlying privacy concerns represent major challenges which are not sufficiently addressed. In this work, we design a cascaded tabular GAN framework (CasTGAN) for generating realistic tabular data with a specific focus on the validity of the output. In this context, validity refers to the the dependency between features that can be found in the real data, but is typically misrepresented by traditional generative models. Our key idea entails that employing a cascaded architecture in which a dedicated generator samples each feature, the synthetic output becomes more representative of the real data. Our experimental results demonstrate that our model well captures the constraints and the correlations between the features of the real data, especially the high dimensional datasets. Furthermore, we evaluate the risk of white-box privacy attacks on our model and subsequently show that applying some perturbations to the auxiliary learners in CasTGAN increases the overall robustness of our model against targeted attacks.
Keywords Generative adversarial networks, tabular data, output validity
1 Introduction
Facilitating information and knowledge sharing within and between organisations is increasingly sought after for attaining growth and development. From a healthcare and medical standpoint, information exchange subsequently contributes to better understanding of diseases and risk factors, more intuitive prognosis by practitioners and effective treatment planning based on previously obtained knowledge. In the financial sector, sharing information between stakeholders leads to improved prediction of corporate bankruptcy and quicker identification of suspicious transaction behaviour that can be potentially linked to organised financial crime. For both fields, sharing the data which contains sensitive patient and client information is subject to GDPR regulations to maintain the confidentiality and privacy of such information. Therefore, institutions are continuously seeking new data anonymisation and synthetic data generation techniques for exchanging domain knowledge without exposing sensitive information.
Ever since their development, generative adversarial networks (GANs) [10] have been increasingly studied for their ability to approximate and model complex data distributions. Despite the early GAN applications being densely focused on the computer vision domain and image generation, GANs are becoming recently researched in other fields such as natural language processing [46] and time-series anomaly detection [24]. In addition, more properties of GANs have emerged such as conditionally generating samples based on a specific target class [26] and generation in conjunction with variational auto-encoders [25].
In contrast, GANs have been significantly less explored for tabular data generation. A tabular dataset typically comprises a mixture of continuous variables and categorical features. Tabular data is common in the medical and the financial domains where fields such age, gender, profession, and income can be commonly found in databases containing numerous records. As opposed to purely numerical data, representing datasets with categorical variables can be particularly difficult in presence of highly-dimensional and strongly correlated features. Furthermore, quantifying the validity of a synthetic tabular dataset can be practically impossible without closely inspecting every generated data sample and deciding whether to accept or reject each examined data record. Notwithstanding, there currently exists no straightforward and unified criteria for evaluating the validity of the output generated by tabular GANs.
To this end, we introduce CasTGAN, which is a generative network framework characterised by multiple generators connected sequentially; each of which is designed to generate a single feature. Meanwhile, a single discriminator validates the output of all the generators while being trained on the output of the final generator in the cascade. In addition, each generator is chained to a corresponding auxiliary learner in order to obtain more insightful losses specific to the individually generated features. This is motivated by the fact that it has been shown that adding more auxiliary classifiers can enhance the quality of the synthetic output images [28]. Therefore, we posit that CasTGAN aims to capture the highly correlated and hierarchical relationship between features, such that the synthetic output produced by our model closely resembles the real data while minimising the inconsistencies in the generated data. This is particularly important for applications where data is widely shared between professionals, and the slightest irregularities in the data can lead to undesired outcomes.
We can thereby summarise our contributions in this work as:
  • Generative architecture: A cascaded based generative framework for producing realistic tabular output which greatly emulates the original data, while significantly reducing the number of invalid synthetic samples.
  • Synthetic data evaluation: A new metric for quantifying the realisticness of the synthetic data when lacking the domain knowledge for the provided data, and extensively evaluate our framework and existing works.
  • Privacy assessment: We launch white-box privacy attacks on our model and analyse how the privacy guarantees and quality of the output are impacted when perturbing the input data during the model training.
The remainder of this paper is structured as follows. In section 2 we present an overview of GANs and the types of GAN privacy attacks, while we further examine the relevant studies in section 3. Section 4 presents a discussion of CasTGAN and a detailed description of our model’s structure. In section 5, we discuss the experimental setup used in this work and the evaluation criteria. We demonstrate our results and a thorough analysis in section 6 and conclude in section 7.
2 Background
2.1 Generative Adversarial Networks
A GAN is characterised by a generator $G$ and a discriminator $D$ playing an adversarial game, where each component attempts to maximise its own benefit [10]. The generator receives a noise input sampled from a random distribution $z\sim p_{z}$ and learns to generate an output in the distribution $x\sim p_{g}$ that matches the structure of the unseen real data $x\sim p_{data}$. Meanwhile, the discriminator has access to the samples produced by the generator and the real data and learns to distinguish between its real and fake inputs. While the output generated by $G$ improves during training as a result of the loss it obtains from the discriminator, the discriminator also becomes increasingly clever in recognising the data produced by the generator. Subsequently, GANs are particularly challenging to train since it must be guaranteed that both the generator and the discriminator maintain their competitiveness without outperforming each other early in the training phase. In the classic GANs, the generator and the discriminator attempt to maximise their objective by minimising the Jenson-Shannon Divergence (JSD), however, using JSD does not guarantee the convergence of losses, hence leading to training instability [10].
The Wasserstein GAN (WGAN) has been proposed as an alternative to the standard GANs in order to augment the training stability in generative models by replacing JSD with the Wasserstein Distance (WD) [1]. The use of the Wasserstein Distance ensures that the model is continuously learning even if the quality of the output is poor, and this is attributed to the smooth gradients produced by the Wasserstein cost function. The initial WGAN relied on weight clipping to enforce the confinement of the discriminator’s weights within a specified range $\left[-c,c\right]$, however, the authors demonstrate that clipping can lead to difficulties with model optimisation [1]. Instead, the use of gradient penalty with WGAN has been proposed to mitigate against the exploding and vanishing gradients of the weights [11]. In this setting, gradients are found using the linear interpolations $\hat{x}\sim p_{\hat{x}}$ between the real and the fake samples, where the distribution of linear interpolation is resembled by $p_{\hat{x}}$. Additionally, the gradient penalty coefficient $\lambda_{GP}$ is used as a parameter for controlling the level at which the gradient penalty affects the discriminator.
The objective function for the WGAN-GP can therefore be represented as:
[…]
$-\lambda_{GP}\>\mathbb{E}_{\widehat{x}\sim p_{\widehat{x}}}\left[\left(\|\nabla_{\widehat{x}}D\left(\widehat{x}\right)\|_{2}-1\right)^{2}\right]$
WGAN-GP is increasingly becoming more prevalent than the classic GANs in applications such as image generation and tabular data generation, as it contributes to more stable learning. In addition, WGAN-GP minimises the effect of mode collapse - that is when the generator learns to "trick" the discriminator by producing a limited number of modes which the discriminator incorrectly classifies as real samples, instead of utilising the entire data feature space.
2.2 Privacy Attacks
In machine learning, membership inference attacks (MIA) aim to identify whether a data sample was used in the training of a machine learning model [33]. For instance, the attackers might try to identify whether the records belonging a client were used for training a loan default prediction model. In this case, the attackers’ objective would be to determine whether the client has taken a bank loan, with such information being used for targeted fraud attempts. Privacy guarantees in machine learning have been extensively studied in the form of analysing the connection to model overfitting [45] and in differential privacy [7].
More recently, MIA have also been explored for generative models. Privacy attacks applied on synthetic samples generated by GAN models aim to reconstruct the real data samples which were used in GAN training. In principle, membership attacks on GANs can be categorised into three types of attacks [12]; full black-box in which attackers have access to only synthetic samples, partial black-box where the attackers have access to the synthetic samples and the latent codes used to generate them, and white-box which assumes that the attackers are able to access the internal parameters of the generator, the discriminator or both. The trade-off between the quality of synthetic samples and the privacy guarantees of GANs have been additionally examined in existing works [41, 15, 5].
3 Related Works
Tabular data is broadly used in regression and classifications tasks, which facilitates a growing interest in tabular data synthesis for machine learning applications, especially in domains with limited training data. Bayesian networks can be used for generating synthetic records by approximating the conditional probablity distribution from the data [47]. While the Bayesian networks can in practice be additionally used for exploring causal relationships between the independent variables, estimating the distributions is often built on simplifying assumptions on the data [9]. Meanwhile, tree-based methods were first utilised for generating partial synthetic data in [31], and has further explored in [40] where adversarial random forest has demonstrated comparable performance to deep learning techniques in terms of synthetic data quality.
Deep neural networks have been widely studied for synthetic data generation, thanks to their capabilities for handling and approximating the distributions of large datasets. Variational auto-encoders (VAEs) [20] estimate the probabilistic distribution of by finding a lower-dimensional latent representation of the data. The application of VAEs has been extended to image data generation [39], oversampling of anomaly event data [14] and tabular data synthesis [42]. An underlying limitation with variational autoencoders is the assumption of the a simple parametric form of the latent space, which leads to a difficulty in capturing complex data distributions [13]. Invertible neural networks have also been proposed for tabular data synthesis through variants based on neural ordinary differential equations [22], copula flows [16] and normalizing flows for private tabular data generation [23].
Within deep learning based generative models, GANs are favourable due to their ability to generate complex synthetic data in an adversarial and unsupervised setting [29]. table-GAN [30] is one of the earliest generative models for producing tabular synthetic output based on adversarial training. Using convolutional neural networks, table-GAN demonstrates that GANs unsurprisingly outperform anonymisation techniques while highlighting the potential privacy risks arising from membership attacks. Meanwhile, in medGAN [6] an autoencoder based generative model is developed for generating high-dimensional medical patient records while shedding light on the privacy risks attributed to the generated data [34]. A long-short term memory (LSTM) architecture for the generator was adopted in [43] demonstrating the potential of recurrent neural networks in synthetic data generation. A GAN approach based on an autoencoder and Kullback-Liebler divergence to tackle mode-collapse in GANs and producing synthetic output comparable to state-of-the-art methods. In CTGAN [42], conditional training of a GAN is carried out by instructing the generator and the discriminator to sample based on randomly selecting feature category in every training iteration, where a highly realistic tabular output can be observed from their evaluation. In [8], the authors propose cWGAN, which is a GAN-based oversampling technique focusing on the generation of samples belonging to the minority class in financial credit datasets. Zhao et. al [48] builds up on existing tabular GAN models to improve the representation of skewed distribution of numerical features of the synthetic output. Finally, Strelcenia et al. comprehensively review the existing tabular GAN literature and evaluate the studied variants for a class imbalance augmentation task [35].
While there is no shortage of novelties in synthetic tabular data generation literature, an underlying challenge remains the proposal of evaluation techniques and criteria for quantifying the reliability and the statistical properties of the synthetic data. A further limitation is the sufficient analysis of data with hierarchical and interdependent variables. Therefore, our focus in this work is proposing a new framework that alleviates the two aforementioned deficiencies in the synthetic tabular data domain.
4 Methodology
Generating synthetic data from unknown and correlated distributions is a non-trivial task. The architecture of CasTGAN is tailored for generating mixed-type features that have similar distributions to the ones observed in a real dataset. Additionally, cascaded structures enable modelling correlations among features in a sequential manner. Given a dataset $X$ with $M$ features, the features of the dataset can be represented as $\{m_{1},m_{2},\ldots,m_{M}\}$.
4.1 Model Architecture
The proposed CasTGAN framework is characterised by a cascade of multiple generators connected sequentially and coupled with auxiliary learners as illustrated in Figure 1.
Figure 1: The model architecture of CasTGAN.
The blocks in Figure 1 denoted with $G_{i}$ and $AL_{i}$, for $i=1,\ldots,M$, are respectively individual generators and auxiliary learners, one per feature $m_{i}$. Each generator $G_{i}$ focuses on generating its target feature using a primary neural network, and is laid out sequentially such that it obtains its inputs from a given noise vector whose components are standard Gaussian and i.i.d., and from the outputs of the previous generator (the only exception being the first generator which only takes a vector of random noise as its input). From a visual perspective, a generic generator $G_{i}$ may then be depicted as in Figure 2.
Figure 2: A close in visualization of one of the various generators in the cascaded layout.
Notation wise, the logic shown in Figure 2 is such that generator $G_{i}$ takes as inputs two objects: the useful outputs coming from $G_{i-1}$ (i.e., the vector $\check{G}_{i}(\phi_{i})$) and the noise vector $z$ (note that the same vector is fed to all the generators as depicted in Figure 1). The generator $G_{i}$ then produces one output, i.e., $G_{i}(\phi_{i})$, that may though be logically split in three distinct components: $Z_{i}$, that will be considered redundant information and that will not be used by the next generator; $\widehat{X}_{i}$, that is the target feature of generator $i$; and $\check{G}_{i-1}(\phi_{i-1})$, that is simply the information from the past generator that will be forwarded to the next one. This means that the output of $G_{i}$ can be formally presented as
$G_{i}(\phi_{i})=\check{G}_{i}(\phi_{i})\oplus Z_{i},$ (2)
while the information that generator $G_{i}$ will pass to $G_{i+1}$ is
$\check{G}_{i}(\phi_{i})=\begin{cases}\widehat{X}_{i}&\text{if }i=1\\ \widehat{X}_{i}\oplus\check{G}_{i-1}(\phi_{i-1})&\text{if }i\geq 2\;.\end{cases}$ (3)
Note that the generator is actually composed by two distinct neural networks: the primary one, whose input is
$\phi_{i}=\begin{cases}z&\text{if }i=1\\ z\oplus\check{G}_{i-1}(\phi_{i-1})&\text{if }i\geq 2,\end{cases}$ (4)
and the secondary neural network, whose input is the noise vector $z$ above and whose output $Z_{i}$ is the redundant information output mentioned above, that will not be passed forward to $G_{i+1}$ but will instead be used by $AL_{i}$.
We note that, as the losses are not backpropagated to the secondary neural network, $G_{i}$ retains its primary objective of generating its target feature based on the input provided to it. Summarizing, the overall cascaded generator structure can be denoted as
$\vec{G}\left(z\right)=\widehat{X},$ (5)
where $\widehat{X}$ is the generated synthetic output.
Based on our literature survey, we observe that some of state-of-the-art tabular GANs employ a conditional setting to enforce the representation of features in both the generator and discriminator [42, 48]. We note that while this is indeed an effective strategy for representing discrete categories and preventing mode collapse, this approach is considerably inefficient for sampling datasets with a large number of categories and few data records, since conditioning on a single random category at every training iteration might not be sufficient to cover all the existing categories in the dataset. In this paper we seek to analyse whether, how much and under which conditions resorting to a series of auxiliary learners – one for each feature – may encourage the models to learn to represent based on the losses traversed, rather than explicitly constraining the model output. The hypothesis is indeed that if multiple auxiliary losses are computed in parallel, the model might be able to improve the learning of categorical interdependence and scale up accordingly to highly dimensional tabular datasets.
4.2 Auxiliary Learners Architecture
GANs for tabular data synthesis are known to be prone to training instability and mode collapse due to the imbalanced feature categories [18]. Conditional GANs [26] have been deployed to generate synthetic output belonging to specific classes. Conditioning both on the generator and the discriminator has been shown to stabilise the training process of GANs.
On the other hand, the use of auxiliary learning for predicting the target variable given the data represents an alternative approach for capturing the characteristics of the data attributed to given target feature. It has been demonstrated that the auxiliary loss further stabilises the training process in comparison to conditional generation, and leads to a representation that it is independent of target label [38]. We observe that while auxiliary learners are traditionally embedded within the discriminator [28], we instead propose designing auxiliary learners as independent structures.
In the CasTGAN, we craft M auxiliary learners $\{AL_{1},\ldots,AL_{M}\}$ for learning to predict the individual features. Due to its scalability on large datasets and the relatively fast convergence speed, we focus on building the auxiliary learners using the Light Gradient Boosting decision trees (LightGBM) [17], which we pre-train prior to the GAN training. An auxiliary learner $AL_{i}$ corresponding to feature $m_{i}$ is trained on $X_{\not\in i}$ in order to predict $X_{i}$. Following standard strategies for such tasks, for predicting the numerical features, the mean-squared error loss is used in the training of the auxiliary learners, whereas cross-entropy loss is used for predicting the categorical and binary variables. As with other decision tree based models, there is no need to one-hot encode the categorical features in $X_{\not\in i}$, but instead the categories are converted into integer encodings. Meanwhile the LightGBM auxiliary learners are capable of handling numerical features with extreme magnitudes, and therefore numerical features are not scaled for auxiliary training.
As LightGBM models have low computational complexity and are generally fast to to train [17], assigning an auxiliary learner for every feature is a reasonable approach for representing the auxiliary loss $L_{AL}$ for predicting a feature given all the other features. It is worth noting that for the early auxiliary learners in the cascaded sequence $AL_{1}\text{ to }AL_{\lceil M/2\rceil}$, the generated data feature space $X_{\not\in i}$ is heavily dominated by redundant variables $Z$ which subsequently lead to increased auxiliary losses. However, these losses help the early generators in producing features that closely match the distributions of the training data. Meanwhile, the task for the later generators and auxiliary learners in the cascade becomes increasingly focused towards generating features that can be predicted from the initially generated target features $\check{G}_{i-1}\left(\phi_{i-1}\right)$.
To ensure that the auxiliary losses $\{L_{AL_{1}},\ldots,L_{AL_{M}}\}$ do not overexceed the generator’s ones, there is the need for scaling down the losses from the auxiliary learners. In this paper we analyze the choice of performing this scaling down by means of constant coefficients $\lambda_{AL}$. In principle, $\lambda_{AL}$ could be a single scalar value that is applied to all the auxiliary losses. However, we consider a vector of auxiliary loss coefficients $\{\lambda_{AL_{1}},\ldots,\lambda_{AL_{M}}\}$ since it is fundamentally important for the early generators to generate variables that conform to the original feature distributions, since this will prompt the next generators in the cascade to effectively learn the feature correlations. Though this a hyperparameter, we set $\lambda_{AL_{1}}=0.75$ and $\lambda_{AL_{M}}=0.10$ while the auxiliary coefficients in between are linearly and equidistantly scaled in the $[0.75,0.10]$ range in our experiments. The same auxiliary setting is applied for all the datasets in this work.
4.3 Training Data Transformation
The main novelty in this paper comes from testing the effects of designing multiple generators in a cascaded layout, where each generator focuses on generating a single data feature. To represent the numerical features we propose to use Variational Gaussian Mixture models (VGM) [3] to estimate the number of modes for a numerical feature, as it has been demonstrated that correctly representing multi-modal numerical data objectively reduces the incidence of mode collapse [34]. In this context, each mode is essentially a Gaussian model on its own, where in the transformation process a single mode is selected for a feature and a scalar value is calculated for quantifying the magnitude of the mode. As such, the transformation of a numerical feature changes the initial unscaled real number into a vector of size equivalent to the number of Gaussian mixture models + 1 (the 1 being the magnitude value of the respective mode and the vector being a one-hot encoded representation of the selected mode). The representation of continuous features using variational Gaussian models is not exclusive to this work as it has been adopted with notable success in earlier tabular GANs [42, 48].
Meanwhile, the categorical features of the training data are transformed into one-hot encodings before being fed to the discriminator. For GANs, the one-hot encoding vectorisation of the categorical features presents an intuitive approach to process the data as it can be scaled and can be appropriately used by the model without issues such as exploding gradients. Furthermore, the use of one-hot encoding simplifies the task of introducing non-linearities by the generator for guaranteeing that the model gradients are differentiable. It is worth reiterating that categorical and numerical transformations of the data for use by the generator and discriminator differ from the representation of the same data used for training and evaluating the auxiliary learners.
4.4 Generators and Discriminator
The generators receive input in the form of noise vector $z$ and the untransformed meaningful output $\check{G}_{i-1}(\phi_{i-1})$. As with other GAN applications, we highlight that using a larger noise vector leads to a better output of the features and can mitigate against mode collapse [34]. Throughout all the experiments, we use a noise vector of size 128, though this is a hyperparameter that can be tuned accordingly [32]. Since each generator dedicates its effort into generating one feature at a time, we use a simple primary neural network of hidden sizes $(128,64)$. Additionally, we use layer normalisation after the hidden dimension [2] for standardizing the weights into zero mean and unit variance and for speeding up the training process. We also use the LeakyReLU activation function with a small negative slope as opposed to ReLU in order to remove the constraints associated with setting the negative gradients to zero.
The dimension of the output layer of the generator for producing the target feature is equivalent to the number of one-hot encodings if the feature is categorical or equal to the number of VGM modes $+1$ if the target feature is numerical. A hyperbolic tangent (tanh) activation is applied to the scalar value of the numerical VGM representation. For the categorical output and the one-hot encoded vector of the VGM vector we use gumbel softmax activations for introducing non-linearities to the output. The Gumbel-softmax works by adding noise from the Gumbel distribution to the vectorised logit output of the generator while maintaining the differentiable nature of the GAN training. The Gumbel-softmax function exhibits also a temperature parameter $\tau$ that may be used to control the diversity of the output generated by the function. In our experiments $\tau=0.8$ was assessed as proper to generate a diversified output that reduces the effects of mode collapse, while conforming to the distribution of categories of the feature within the training set.
We then note the risk that the discriminator may learn to distinguish between real and generated data by discriminating between the hard one-hot encoded real data and the float values from the generator. To minimize this risk we add an i.i.d. Gaussian noise distributed as $\mathcal{N}(0,0.01)$ [1] to the columns of the real samples before feeding them to the discriminator. Consequently, all the inputs to the discriminator (i.e., numerical and categorical features of real and generated data) are float values. The weights of the discriminator are then trained using only the outputs from the final generator $G_{M}$. The parameters of the various generators are though updated based on the loss of the discriminator, that is thus computed for this reason.
We maintain a simple architecture for the discriminator comprising two hidden layers of sizes $(256,128)$. As with the generator, layer normalization and LeakyReLU activations are used between the hidden layers. The final layer consists of a single output node without an activation function. To alleviate against mode collapse & GAN training instability issues we additionally compute the Wasserstein loss [1] with gradient-penalty [11] for the calculation of the discriminator losses.
As the CasTGAN is built up using $M$ generators, we have multiple min-max games between the generators and the discriminator. Therefore, the value function for the discriminator can be expressed as
[…]
$-\>\mathbb{E}_{z\sim p_{z}}\left[D\left(\vec{G}(z)\right)\right]$
$-\>\lambda_{GP}\mathbb{E}_{\widehat{x}\sim p_{\widehat{x}}}\left[\left(\|\nabla_{\widehat{x}}D\left(\widehat{x}\right)\|_{2}-1\right)^{2}\right]$
and value function for generator $G_{i}$ is hence given by
[…]
$+\>\lambda_{AL_{i}}\>\mathbb{E}_{\phi\sim p_{\phi}}\left[L_{AL_{i}}\right].$
5 Experimental Setup
Evaluating the performance of GANs is a non-trivial task, and this can be evident from the variety of GAN evaluation criteria in literature. In this section, we describe the experimental design for CasTGAN, with a thorough discussion of the metrics used for the model’s analysis. We implemented our model using PyTorch in Python on a Linux Ubuntu 20.04 machine running on AMD Ryzen Threadripper 3990X and Nvidia GeForce RTX 3090. Our CasTGAN source code is publicly available [^1] https://github.com/abedshantti/CasTGAN.
5.1 Datasets
Our CasTGAN is designed for synthesis of tabular data that can be typically found in the financial and healthcare sectors. We therefore use tabular mixed datasets characterised by a combination of categorical and numerical features. We use four datasets where the task is the binary classification of the target label - Adult [21], Bank Marketing [27], Taiwan Credit [44] and Diabetes [36]. Meanwhile, we use the House Prices [19] and Cars [37] datasets for regression. We additionally highlight that binary columns in the datasets are handled as categorical variables. An overview of the datasets used is presented in Table 1. For synthesising data with CasTGAN and other baselines, we use 50% of the datasets’ total number of samples for training the models. The remaining 50% of the data samples is dedicated for evaluating the generated synthetic output.
Table 1: Datasets used in this study.
Dataset Samples Num. Cat. Unique Task
features features categories
Adult 32561 6 9 104 Binary classification
Bank 45211 10 11 55 Binary classification
Credit 30000 14 10 79 Binary classification
Diabetes 253680 7 15 30 Binary classification
Cars 472336 5 8 2409 Regression
Housing 21613 17 2 72 Regression
5.2 Baselines
We compare the synthetic output of CasTGAN against five state-of-the-art tabular generative adversarial network models: table-GAN [30], medGAN [6], CTGAN [43], cWGAN [8] and CTAB-GAN [48]. Given that some of the datasets that we use in this study were not evaluated previously by the existing methods, we selected the optimal hyperparameters recommended by the baseline methods’ authors in this study.
5.3 Evaluation Criteria
5.3.1 Train on Synthetic, Test on Real (TSTR)
We measure reliability of the synthetic output produced by CasTGAN by training machine learning models on the generated data. We fitted three machine learning models on the generated output - namely AdaBoost, random forest and logistic regression for classification tasks and AdaBoost, decision trees and Linear SVM for regression tasks. We then used the trained models to predict the target label of the test data and we report the precision-recall area under curve (PR-AUC) for binary classification tasks and the R2 score for regression tasks. Since our main objective is to measure the machine learning utility of the generated data rather than assessing the individual performance of each classifier, we average the metrics produced by the three machine learning models.
5.3.2 Univariate Distributions
We also assess how well do the individual features generated by CasTGAN resemble the features of the real data. As we quantitatively analyse how well our model learns the univariate feature distributions, we first one-hot encode and normalise the synthetic and real datasets. We calculate the dimension-wise mean of the individual features of the synthetic output and training data and report the RMSE between the real and the synthetic output. Additionally, we report the Kolmogorov-Smirnov two-sample test score [4] between the real univariate features and the synthetic ones.
5.3.3 Correlation and Diversity
Measuring the validity of the synthetic output represents a significant challenge for GAN frameworks. In computer vision applications of GANs, the synthetic images can be in some cases distinguished from the real images by a human observation of irregularities in the output such as pupil orientation in human eyes and out-of-place pixels. In tabular GANs, similar challenges exist as there is no standard approach in the literature for quantifying the proportion of invalid samples in the synthetic output. For instance, a common example in the tabular synthesis literature is highlighting how an entry such as gender = "Female" and diagnosis = "Prostate Cancer" is by definition an invalid record, as no such entry exists in the original data, nor can a female be diagnosed with prostate cancer by a physician. In financial datasets, such nested relationships also exist between features if for example we inspect a dataset with a city column and a country column. In such cases, a record with city = "Buenos Aires" and country = "Malta" is by definition invalid. Meanwhile, entries with city = "Alexandria" and country = "United States" is valid as Alexandria exists in the United States even though it is more commonly attributed to the country "Egypt". It is for this reason that quantifying the invalid output generated by tabular GANs is no easy task, even in the presence of domain knowledge.
While a GAN model needs to ensure that its synthetic output is as valid as possible, there also needs to be some considerations for the diversity of the generated output. As such, the generative models should not significantly restrict the possible feature combinations between the different categorical features. Ensuring the diversity of the synthetic output enables the model to be less deterministic and increases its robustness against privacy attacks that aim to identify sensitive information. Therefore, the GAN model should be encouraged to explore unique feature combinations as long as such combinations can be considered valid.
Given that public datasets are used in this work for the purpose of reproducability, we do not have the full domain knowledge for these datasets, thus, we propose an alternative method for quantifying the validity of the synthetic data. First, we consider calculating the difference in feature correlations between the training data and the fake data. For computing the correlations between numerical features we use the Pearson’s correlation coefficient, while the Cramer’s V measure is used for capturing the correlation between categorical features. The correlation score is found by calculating the RMSE score between the elements of the triangular matrix of the synthetic dataset and the real dataset.
For measuring the diversity of the categorical output, we propose a new metric - Unique Pairwise Categorical Combinations (UPCC). In essence, we count the total number of unique interactions between any pair of categorical features in the dataset. For instance, in the Adults dataset the combinations [education = "Bachelors" and marital-status = "Never-married"], [education = "Bachelors" and sex = "Male"] and [marital-status = "Never-married" and sex = "Male"] each counts as a single pairwise combination, regardless of how many times they appear in the data. A reliable model therefore ensures that the UPCC of its output should be comparable to that of the original data. Subsequently, the UPCC Ratio is the number of unique pairwise combinations of synthetic output divided by the total number of unique combinations of the training data.
Finally, we estimate the validity of the model’s output by dividing the correlation RMSE score of the model by the UPCC Ratio. We name this measure as the CORDV score. A lower CORDV score indicates that the model is able to minimise the difference in feature correlation between its synthetic output and the real data, while simultaneously not impeding its ability in generating unique feature combinations. Meanwhile, a worse generative model is reflected a greater CORDV score indicating that the model poorly captures the correlations while potentially restricting the uniqueness of the categorical pairs.
5.3.4 White-box privacy attacks
Traditionally, white-box membership inference attacks on GANs assume that the attacker has access to the synthetic data and at least one generative component of the model. In this work, we formulate white-box privacy attacks in a different setting. We highlight that while using multiple auxiliary learners help in generating more realistic and reliable synthetic output, the use of multiple auxiliary learners leads to a more susceptible model for privacy breaches by attackers.
In this work, we devise white-box attacks by assuming that an attacker has access to the trained auxiliary learners and attempts to reconstruct training samples through an iterative process of estimating a hidden feature. In essence, the attacker with the synthetic data will at a given time remove one column from the data, use the corresponding the auxiliary learner to predict the masked feature using the remaining features, and then replacing the masked column with the predicted output from auxiliary learners. In this setting, a single iteration refers to a walk-through over all the auxiliary learners and subsequently replacing all the columns in the dataset once.
For evaluating how effective such white-box attacks on our model, we control the training the of the auxiliary learners using a perturbation parameter $\epsilon$. The perturbation parameter translates to the proportion of label samples that are modified when training the auxiliary learners prior to the GAN training. For an auxiliary learner corresponding to a numerical column $X_{i}$, we perturb the numerical variables such that perturbed variable for a given sample $\tilde{x}$ can be calculated as $\tilde{x}_{i}=x_{i}+\alpha x_{i}$, where $\alpha$ is a floating number randomly sampled from $[-1.0,1.0]$. Meanwhile, we perturb the categorical features by randomly selecting a category from the list of all the unique categories of the said feature. In our analysis, we experiment with $\epsilon=0.0$, implying that no perturbation takes place, and gradually increment this value to $\epsilon=0.3$, implying that 30% randomly chosen samples for each auxiliary learner were perturbed prior to the auxiliary training.
Furthermore, we analyse whether an attacker possessing the original data preprocessing transformers has an additional advantage in recovering the training samples. We hypothesise that an attacker with access to the data transformations used for the auxiliary training can simply convert the synthetic to a data structure that aligns with the existing transformer. On the other hand, an attacker without access to the preprocessors needs to fit and transform the data independently before launching membership attacks. This is especially prevalent for categorical features, where transforming the categories into ordinal encodes that do not match the ones learned by the auxiliary learners might lead to less effective attacks.
6 Results
6.1 Machine Learning Utility
We use the synthetic data generated by CasTGAN for fitting machine learning classification and regression supervised models on the six datasets and compare our performance on the test set against models fitted on the training set and models fitted on the five synthetic output of the five baseline methods. The results are computed in Table 2.
Table 2: Binary classification (PR-AUC score) and regression (R2 score) evaluation on the test sets.
Adult Bank Credit Diabetes Cars Housing
Identity 0.7744 0.6085 0.5374 0.3980 0.7604 -0.3696
table-GAN 0.2225 0.0954 0.2084 0.1585 -2.0148 -83.1326
medGAN 0.5777 0.2729 0.3050 0.3241 0.7491 -0.3845
CTGAN 0.6932 0.4805 0.4558 0.3711 0.5327 -0.6533
cWGAN 0.3030 0.2692 0.2166 0.3325 -319.0963 -1.1054
CTAB-GAN 0.7192 0.4920 0.4911 0.3589 - -0.7284
CasTGAN 0.6718 0.5657 0.4995 0.3866 0.5566 -0.4433
From Table 2 we can observe that the TSTR metrics for our CasTGAN is consistently within the best performing synthetic output, outperforming all the baselines on three out of six datasets. Meanwhile, we note that medGAN demonstrated the best results on regression datasets. In addition, CTGAN and CTAB-GAN have comparatively close performance to our method and the training examples. We note that we were unable to run the highly-dimensional Cars dataset on CTAB-GAN to exceedingly large memory requirements attributed to representing wide datasets in convolutional neural network GAN-based approaches.
6.2 Univariate Similarity
To quantitatively analyse how well CasTGAN represents the feature distributions of the original data, we compare the Euclidean distance RMSE and Kolmogorov-Smirnov statistic between the synthetic data and the real data in Table 3. Generally, we note that our CasTGAN, medGAN and CTAB-GAN dominate the dimension-wise statistical similarity test. It can be observed that CasTGAN represents the features of the Adult and Cars datasets particularly well, while performing comparatively on the remaining datasets. We can therefore deduce that CasTGAN is particularly useful for datasets with a greater number of unique categories.
Table 3: Univariate dimension-wise statistical comparison.
(a) Adult, Bank and Credit datasets
Adult Bank Credit
Euc. RMSE KS statistic Euc. RMSE KS statistic Euc. RMSE KS statistic
table-GAN 0.2249 0.6499 0.3485 0.5969 0.3218 0.8794
medGAN 0.0599 0.1290 0.0636 0.1963 0.0570 0.3038
CTGAN 0.0377 0.1502 0.0415 0.1076 0.0786 0.1715
cWGAN 0.0406 0.1111 0.0616 0.5536 0.0167 0.1906
CTAB-GAN 0.0186 0.0812 0.0396 0.0871 0.0335 0.0959
CasTGAN 0.0111 0.0908 0.0276 0.1470 0.0283 0.1327
(b) Diabetes, Cars and Housing datasets
Diabetes Cars Housing
Euc. RMSE KS statistic Euc. RMSE KS statistic Euc. RMSE KS statistic
table-GAN 0.4087 0.4247 0.0531 0.7088 0.1476 0.6606
medGAN 0.0188 0.0447 0.0124 0.1323 0.0141 0.1812
CTGAN 0.0767 0.1024 0.0067 0.0785 0.0571 0.1682
cWGAN 0.1133 0.2809 0.1065 0.2890 0.1096 0.2956
CTAB-GAN 0.0648 0.0798 - - 0.0150 0.0989
CasTGAN 0.0834 0.0780 0.0019 0.0457 0.0184 0.1651
6.3 Output Validity
We quantify the validity of the output by considering the number of unique pairwise categorical combinations (UPCC), the correlation error between the synthetic and the real data, and the CORDV score, which is essentially the correlation divided by the UPCC ratio. From Table 4, we can observe how the different generative methods rank among the three aforementioned metrics. First, we note that CTGAN and medGAN perform well in generating a large number of unique feature combination, in most cases, even more than the number of combinations that can be found in the training data. This is particularly impressive for the Cars dataset, where both models managed to generate more than twice the number of categorical combinations of the training set. Similarly, we also observe that CTAB-GAN performs relatively well in exploring diverse categorical combinations. Meanwhile, we notice that our CasTGAN is more conservative when it comes to generate unique categorical combinations. For all the dataset, CasTGAN produces a marginally lower number of pairwise combinations than can be typically found in the training set. Moreover, the UPCC can be a good indicator of mode collapse and this is reflected by the significantly low UPCC values for table-gan and cWGAN, where it can be deduced that these models generated a limited number of modes for some categories.
Table 4: Diversity and correlation comparison.
(a) Adult, Bank and Credit datasets
Adult Bank Credit
UPCC Corr. RMSE CORDV UPCC Corr. RMSE CORDV UPCC Corr. RMSE CORDV
Identity 3004 1244 1426
table-GAN 1168 0.2010 0.5170 461 0.2516 0.6789 878 0.3001 0.4874
medGAN 3769 0.1708 0.1361 1233 0.1371 0.1383 1971 0.1826 0.1321
CTGAN 3435 0.0570 0.0498 1306 0.0749 0.0713 2501 0.1003 0.0572
cWGAN 2404 0.0718 0.0897 1255 0.1128 0.1118 1963 0.0804 0.0584
CasTGAN 2369 0.0368 0.0466 1163 0.0822 0.0879 1191 0.1130 0.1353
(b) Diabetes, Cars and Housing datasets
Diabetes Cars Housing
UPCC Corr. RMSE CORDV UPCC Corr. RMSE CORDV UPCC Corr. RMSE CORDV
Identity 420 62872 85
table-GAN 134 0.1482 0.4646 28 0.4015 901.5202 62 0.3493 0.7890
medGAN 3769 0.1708 0.1361 1233 0.1371 0.1383 1971 0.1826 0.1321
CTGAN 420 0.0575 0.0575 128405 0.1677 0.0821 140 0.1755 0.1065
cWGAN 420 0.0512 0.0512 - - - 107 0.1384 0.1100
CasTGAN 406 0.0490 0.0507 31782 0.1351 0.2672 83 0.0880 0.0909
In contrast, it can be observed from Table 4 that CasTGAN generally outperforms the other baselines in capturing the feature correlations of the datasets. The lower correlation RMSE score entails that CasTGAN prioritises the representation of correlations and feature interdependence in the real data. Meanwhile, the CORDV score aims to quantify the trade-off between the diversity and the proximity of the synthetic data to the real data. We observe from the results that the highest CORDV scores are evenly split among our CasTGAN and CTGAN.
Despite lacking the full domain knowledge in our datasets, we nonetheless measure the synthetic invalidity on the Adult, Cars and Housing datasets as follows:
  • Adult dataset: we use the fields "relationship" and "sex" to calculate the number of invalid records. We posit that if "relationship" = "Husband", then the sex feature needs to be set to "Male". Likewise, the "relationship" = "Wife" needs to align the gender field assigned as "Female". This is not based on our assumptions, but rather running explanatory data analysis on the training set confirms that the records are matched in such a manner.
  • Cars dataset: we use the fields "make" and "model" and classify a synthetic sample as invalid if the synthetic car’s "model" does not in fact belong to the "make" that can be found in training data.
  • Housing dataset: the housing dataset consists of the fields "year built" and "year renovation". Logically, a property cannot be renovated before it was built, and further inspecting the data indeed confirms that there are no observations with "year renovation" before "year built".
Based on the aforementioned fronts, the ratio of invalid records generated by CasTGAN and three baseline methods are demonstrated in Table 5. We chose not to include table-GAN and cWGAN in the comparison as the synthetic output is characterised by major mode collapse.
Table 5: The ratio of invalid synthetic records.
Adult Cars Housing
medGAN 0.2905 0.9111 0.5929
CTGAN 0.0982 0.4289 0.5475
CTAB-GAN 0.0414 - 0.4442
CasTGAN 0.0067 0.2782 0.2298
From Table 5 we can observe that CasTGAN remarkably reduced the number of invalid synthetic records of the Adult dataset. Our method also significantly decreased the number of invalid records in the Cars dataset. This resembles a major improvement from CTGAN, while noting the challenging nature of modelling the Cars dataset due to the large number of categories present. Furthermore, it is evident that CasTGAN outperforms the other generative approaches on the numerical features of the Housing dataset.
6.4 Robustness Against Privacy Attacks
For conducting white-box privacy attacks, we set the number of attacking iterations to five, where each feature in the synthetic data is updated five times based on the output of the auxiliary learners. Moreover, the membership attacks are launched on 10% of the total number of overall samples. We highlight that the ratio of attacked samples does not impact the evaluation of the robustness of our approach as we only compute the attack distance metrics with respect to the attacked samples. The Euclidean distance of the attacked samples to the training data and to the synthetic data prior to the membership attacking is computed in Table 6.
Table 6: White box privacy attacks on the auxiliary learners proximity measures.
Access to only AL Access to AL and Preprocessing
Euc. to train Euc. to pre-attack syn Euc. to train Euc. to pre-attack syn
Adult $\epsilon=0.0$ 1.2202 1.4437 0.1256 0.3652
$\epsilon=0.1$ 1.4250 1.7289 0.2025 0.7348
$\epsilon=0.2$ 1.4747 1.8288 0.6848 1.4039
$\epsilon=0.3$ 2.4468 2.5217 0.7413 1.2509
Bank $\epsilon=0.0$ 2.2275 2.4975 1.2225 1.7366
$\epsilon=0.1$ 3.0541 3.0272 2.4020 2.4182
$\epsilon=0.2$ 5.6174 5.9515 5.3565 5.8131
$\epsilon=0.3$ 6.7180 7.1127 6.7095 7.0654
Credit $\epsilon=0.0$ 0.3740 0.6550 0.1157 0.3179
$\epsilon=0.1$ 0.6196 0.7916 0.1094 0.2804
$\epsilon=0.2$ 0.3918 0.5797 0.0732 0.2773
$\epsilon=0.3$ 1.5533 1.7834 0.6349 0.8470
Housing $\epsilon=0.0$ 0.3696 0.4905 0.2026 0.4132
$\epsilon=0.1$ 4.5976 4.7431 4.4731 4.5013
$\epsilon=0.2$ 16.2106 16.6204 16.1899 16.5996
$\epsilon=0.3$ 28.6935 28.8148 28.6990 28.8205
From Table 6, it can be observed that the perturbation coefficient $\epsilon$ greatly impacts the closeness of attacked samples to the training data. For unperturbed and minimally perturbed data features, it can be noticed that the attacked samples are relatively close to the training samples, indicating that the attackers might succeed in recovering training datapoints. We observe that the proximity to the training samples increases for greater $\epsilon$ values, which demonstrates the additional privacy guarantees that can be provided when altering the labels. We additionally notice how access to the data processors gives a major advantage to the attackers attempting to recover the training samples. This holds true especially for the Adult dataset, where using the trained label encodings of the auxiliary learners lead to more targeted attacks that are even closer to the training samples than attackers on unperturbed data with no access to the data preprocessors. Another interesting observation is that the attacks on the Housing dataset are greatly impacted by incremented in $\epsilon$, which is plausible, given that the dataset mainly consists of numerical features.
In addition to the proximity to the training samples, we also investigate whether perturbing the labels of the auxiliary learners can contribute to a reduction in the quality of unattacked synthetic data as demonstrated in Table 7. For the Adult, Bank and Credit datasets it can be evident that applying perturbations insignificantly impacts the the evaluation metrics of the synthetic datasets. We observe that the PR-AUC scores on the test data and KS statistic for univariate distributions are minimally influenced by the changes in $\epsilon$. In contrast, it appears that perturbing the data impacts the CODRV scores as a result of the correlation errors between the synthetic and the real datasets. Interestingly, applying perturbations on the Housing dataset appear to improve the quality of the synthetic output in addition to increasing its robustness against white-box privacy attacks.
Table 7: Impact of auxiliary learners imputation on the synthetic output.
TSTR KS CODRV
Adult $\epsilon=0.0$ 0.6718 0.0908 0.0467
$\epsilon=0.1$ 0.6683 0.1173 0.0380
$\epsilon=0.2$ 0.6731 0.1560 0.0393
$\epsilon=0.3$ 0.6647 0.1346 0.0425
Bank $\epsilon=0.0$ 0.5657 0.1470 0.0879
$\epsilon=0.1$ 0.5384 0.1020 0.0847
$\epsilon=0.2$ 0.5551 0.1253 0.0868
$\epsilon=0.3$ 0.5475 0.1130 0.0960
Credit $\epsilon=0.0$ 0.4995 0.1327 0.1353
$\epsilon=0.1$ 0.5140 0.1069 0.1504
$\epsilon=0.2$ 0.4722 0.1170 0.1434
$\epsilon=0.3$ 0.5095 0.1103 0.1822
Housing $\epsilon=0.0$0.4430 0.1651 0.0909
$\epsilon=0.1$0.5602 0.2185 0.0864
$\epsilon=0.2$0.3257 0.1554 0.0961
$\epsilon=0.3$0.2966 0.1381 0.0828
7 Conclusion
In this work, we presented CasTGAN as a generative framework for creating synthetic tabular data samples that are representative of the real data attributes. Our motivation for this work stems from the need for realistic tabular data that can be exchanged amongst experts, while focusing on the reliability and the sensitivity of such information. We therefore directed our focus towards generating fake output that capture the correlations and interdependence between the data features. We demonstrated that our cascaded generator architecture supported by auxiliary learners are able to generate realistic output given highly dimensional and largely imbalanced tabular datasets. Our results indicate that CasTGAN is capable of significantly reducing the number of invalid records while exhibiting strong statistical and correlational similarities to the real data. We further evaluated the robustness of our model against targeted privacy attacks and showed that perturbing the auxiliary learners by a small scale can mitigate against attacks aiming to recover the real data samples.
Given the challenging nature of generating realistic synthetic tabular data, there are several paths for future work. This work can be extended by incorporating additional data types within the tabular data generation such as free text and timestamps. Moreover, we intend to explore how our approach can generate more diversified combination of categories, while maintaining its ability in minimising the number of invalid data records.
Acknowledgments
This project was supported by DNB ASA through the funding of this research.
Appendix A Training Details
A.1 Hyperparameters
Table 8: Hyperparameters used for all datasets
Hyperparameter Value
Epochs 300
Batch size 512
Generators noise input dimension 128
Generators noise input distribution $\sim\mathcal{N}(0,1)$
Generators primary networks hidden sizes [128, 64]
Generators primary networks inter-hidden layer activations LeakyReLU
Generators primary networks activation normalizer Layer Normalization
Generators primary networks numerical feature activation layer tanh
Generators primary networks categorical feature activation layer gumbel softmax$(\tau=0.8)$
Generators secondary networks hidden sizes [32]
Generator optimizer Adam($\alpha$ = 2 $\times$ 10-4, $\beta_{1}$ = 0.50, $\beta_{2}$ = 0.99)
Discriminator primary networks hidden sizes [256, 128]
Discriminator inter-hidden layer activations ReLU
Discriminator activation normalizer Layer Normalization
Discriminator optimizer Adam($\alpha$ = 2 $\times$ 10-4, $\beta_{1}$ = 0.50, $\beta_{2}$ = 0.99)
Discriminator real input noise perturbation $\sim\mathcal{N}(0,0.01)$
Discriminator updates per generators update 1
Auxiliary learners boosting iterations 150
Auxiliary learners number of leaves 31
Auxiliary learners learning rate 0.10
Auxiliary learners early stopping 10 rounds
Wasserstein loss gradient penalty $\lambda_{GP}$ 10
Auxiliary learner 1 loss coefficient $\lambda_{AL_{1}}$ 0.75
Auxiliary learner $M$ loss coefficient $\lambda_{AL_{M}}$ 0.10
A.2 Losses
(a) Adult Dataset Losses
(b) Bank Dataset Losses
(c) Cars Dataset Losses
(d) Credit Dataset Losses
(e) Diabetes Dataset Losses
(f) Housing Dataset Losses
Figure 3: Training losses
A.3 Training Elapsed Duration
Figure 4: Comparison of run times between the different baselines
Appendix B Further Results
B.1 TSTR
Table 9: Adult Dataset Classification
Table 10: Bank Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
train 0.8550 0.9075 0.6663 0.7744
tablegan 0.5617 0.4102 0.1376 0.2225
medgan 0.7920 0.8038 0.4888 0.5777
ctgan 0.8304 0.8751 0.6235 0.6932
cWGAN 0.7586 0.5887 0.0002 0.3030
ctab-gan 0.8400 0.8849 0.6542 0.7192
CasTGAN 0.8272 0.8779 0.6380 0.6718
Accuracy ROCAUC F1-score PR-AUC
train 0.9084 0.9342 0.4978 0.6085
tablegan 0.2149 0.3910 0.1670 0.0954
medgan 0.8803 0.6492 0.1547 0.2729
ctgan 0.8979 0.8474 0.3492 0.4805
cWGAN 0.8337 0.6864 0.1297 0.2692
ctab-gan 0.8970 0.8768 0.3550 0.4920
CasTGAN 0.8991 0.9187 0.5823 0.5657
Table 10: Bank Dataset Classification
Table 11: Credit Dataset Classification
Table 12: Diabetes Dataset Classification
Accuracy ROCAUC F1-score PR-AUC
train 0.8201 0.7685 0.4663 0.5374
tablegan 0.2185 0.4605 0.3482 0.2084
medgan 0.6807 0.6146 0.3448 0.3050
ctgan 0.7826 0.6845 0.4561 0.4558
cWGAN 0.7760 0.5253 0.0053 0.2166
ctab-gan 0.8138 0.7136 0.4708 0.4911
CasTGAN 0.8048 0.7315 0.2845 0.4995
Accuracy ROCAUC F1-score PR-AUC
train 0.8623 0.8142 0.2545 0.3980
tablegan 0.8604 0.5222 0.0000 0.1585
medgan 0.8519 0.7544 0.2473 0.3241
ctgan 0.8333 0.7990 0.3924 0.3711
cWGAN 0.8443 0.7636 0.2749 0.3325
ctab-gan 0.8432 0.7857 0.3519 0.3589
CasTGAN 0.8557 0.8077 0.3336 0.3866
Table 12: Diabetes Dataset Classification
Table 13: Cars Dataset Regression
Table 14: Housing Dataset Regression
RMSE R2 Score
train 4330.3428 0.7604
tablegan 16737.5294 -2.0148
medgan 4822.8477 0.7491
ctgan 6558.5413 0.5327
cWGAN 166818.8351 -319.0963
CasTGAN 6396.3703 0.5566
RMSE R2 Score
train 393507.3597 -0.3696
tablegan 2943381.0066 -83.1326
medgan 405452.4482 -0.3845
ctgan 458694.4080 -0.6533
cWGAN 518963.3807 -1.1054
ctab-gan 472485.7269 -0.7284
CasTGAN 419848.1947 -0.4433
Table 14: Housing Dataset Regression
B.2 Statistical Similarity
(a) Categorical features
(b) Numerical features
Figure 5: Adult dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 6: Bank dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 7: Cars dataset statistical similarity
(a) Categorical features
(b) Numerical features
Figure 8: Diabetes dataset statistical similarity